OpenStudy (anonymous):

when n->infinit, the limit of (1/n^3)(1^2+3^2+...+(2n-1)^2) =?

OpenStudy (anonymous):

is (1^2+3^2+...+(2n-1)^2) in top or bottom?

OpenStudy (anonymous):

top

OpenStudy (anonymous):

(1^2+3^2+...+(2n-1)^2)/n^3 =? this looks better.

OpenStudy (anonymous):

\[\lim_{n\to \infty } \, \frac{1^2+3^2+\text{...}+(2n-1)^2}{n^3}\]

OpenStudy (anonymous):

yes . You use the tools to type this ?

OpenStudy (anonymous):

yep, latex

OpenStudy (anonymous):

Is latex the name of the tool? How can i get it? Can you give me a link?

OpenStudy (anonymous):

no , you type the code, for example, I type this :\lim_{n\to \infty } \, \frac{1^2+3^2+\text{...}+(2n-1)^2}{n^3}

OpenStudy (anonymous):

between \[ and

OpenStudy (anonymous):

Wow..

OpenStudy (anonymous):

That is right. but the result is still series

OpenStudy (anonymous):

I am trying by squeez theorem..

OpenStudy (anonymous):

The answer from the textbook is 4/3.

OpenStudy (nilankshi):

help me also pls

OpenStudy (anonymous):

My summation was wrong, \[\lim_{n\to \infty } \, \frac{\sum _{k=1}^n (-1+2 k)^2}{n^3}\] \[\lim_{n\to \infty } \, \frac{\sum _{k=1}^n 4 (-1+2 k)}{3n^2}\] \[\lim_{n\to \infty } \, \frac{\sum _{k=1}^n 8}{6}\] \[\lim_{n\to \infty } \, \frac{8\sum _{k=1}^n 1}{6}\]

OpenStudy (anonymous):

ignore my last two step : \[\lim_{n\to \infty } \, \frac{\sum _{k=1}^n 8}{6n}\] \[\lim_{n\to \infty } \, \frac{8\sum _{k=1}^n 1}{6n}\] \[\lim_{n\to \infty } \, \frac{8n}{6n}\] \[\lim_{n\to \infty } \, \frac{8n}{6n}=\frac{4}{3}\]

OpenStudy (anonymous):

I see... Thank you so much . I learn a new skill for using L'Hospital's Rule. Thank you

OpenStudy (anonymous):

you are welcome

OpenStudy (anonymous):

I don't know the L'Hospital's Rule can use in that way before. I am so happy to learn it...HAHA

OpenStudy (anonymous):

yeah they are usually best and easy way to determine limit. To do l'hospital , we need inderterminant of either 0/0 or infinity/infinity

OpenStudy (anonymous):

Use l'hospital directly in the \[\sum_{?}^{?}\] is wonderful. I learn the skill.

OpenStudy (anonymous):

Could you tell me more about the way you use l'hospital: I mean the way you take derivative directly in the ∑ .

OpenStudy (anonymous):

I have another example: Prove that : \[if:\mathop {\lim }\limits_{n \to \infty } {a_n} = a\] then \[\mathop {\lim }\limits_{n \to \infty } \frac{{{a_1} + 2{a_2} + ... + n{a_n}}} {{{n^2}}} = \frac{a} {2}\]

OpenStudy (anonymous):

Can i use l'hospital rule like you did: \[\mathop {\lim }\limits_{n \to \infty } \frac{{\sum\nolimits_{k = 1}^n {k{a_k}} }} {{{n^2}}} = \mathop {\lim }\limits_{n \to \infty } \frac{{\sum\nolimits_{k = 1}^n {{a_k}} }} {{2n}} = \mathop {\lim }\limits_{n \to \infty } \frac{{an}} {{2n}} = \frac{a} {2}\] Am i right? The new way of using l'hosptial rule is a little confused.